OpenStudy (anonymous):

The sum of two numbers is 205. One number is 10 more than twice the other. Find the larger of the two numbers.

OpenStudy (callisto):

let x be the the smaller number. the larger number = 2x+10 since sum of 2 numbers =205 x + (2x+10) =205 3x = 205-10 x= (195)/3 =65 So the larger number = 205 - 65 = work it out yourself !

OpenStudy (anonymous):

x + y = 205 x = 2y + 10 Solve for y, substituting in x 3y = 195 y = 65, so x = ?

OpenStudy (anonymous):

75??

OpenStudy (anonymous):

Re-read the question. Look at the equations, Does x= 75 satisfy them? (it does not)

OpenStudy (anonymous):

140

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

because 145 + 65 =205 correct??

OpenStudy (callisto):

nope, 145+65 = 210

OpenStudy (anonymous):

oops i meant 140 lol